# Eulerian and Lagrangian electron energization during magnetic reconnection

### v1

@@ 3.3 Eulerian vs Lagrangian energization
On system scales, however, the picture is very different. For our choice of boundary conditions, the system is closed in the sense that there is no net particle energy flux into or out of the simulation domain. Therefore, when we integrate the divergence terms of Eq. (4) over the whole simulation, the corresponding surface terms vanish, leaving us with $\boldsymbol{J}_{e}\cdot\boldsymbol{E}\approx\partial\mathcal{E}_{\text{EFM}}/\partial t$, as seen in Figure 5i. This is also why the LGCM can more accurately reproduce $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ integrated over the entire simulation domain, when $\mu$-conservation applies (Figure 2e). In that case, Eq. (2) describes the evolution of the total energy of all guiding centers, and since they are confined to the simulation domain, this is the same thing as the evolution of the total electron energy in the simulation domain. So, when we look at a closed system in its entirety, indeed $\int d^{3}x(d\mathcal{E}_{\text{LGCM}}/dt)=\int d^{3}x(\partial\mathcal{E}_{\text{EFM}}/\partial t)=\int d^{3}x\boldsymbol{J}_{e}\cdot\boldsymbol{E}$.
Until this point, we have followed TenBarge et al., (2024) and uncritically assumed that, since $d\mathcal{E}_{\text{LGCM}}/dt$ accurately describes the energization of individual guiding centers, it must equal the work done on the electrons as quantified by $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. In other words, since we can write Eq. (2) as the scalar product of a net current density $\boldsymbol{J}_{\text{e,LGCM}}$ and $\boldsymbol{E}$, $\boldsymbol{J}_{\text{e,LGCM}}$ should be equal to the bulk fluid current density $\boldsymbol{J}_{e}$. This, however, is evidently not the case. As already mentioned by TenBarge et al., (2024), the guiding center model is a particle model that is gyroperiod integrated, and it consequently does not describe certain bulk fluid properties such as the diamagnetic drift and the corresponding currents. Therefore, in general $\boldsymbol{J}_{e,LGCM}\neq\boldsymbol{J}_{e}$, since $\boldsymbol{J}_{e}$ describes the total electron current. In hindsight, it is therefore quite clear that we should never have expected $d\mathcal{E}_{\text{LGCM}}/dt$ to accurately model $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ to begin with, even when $\mu$ is conserved.
@@ 4.1 Interpretation of EFM energization terms
To illustrate how the above scenario affects the local energy evolution of the plasma, we consider an Eulerian volume element that is initially in the high-pressure region. At the start the uniform plasma is $\boldsymbol{E}\times\boldsymbol{B}$-drifting across the volume element, resulting in $\nabla\cdot\boldsymbol{Q}=\partial\mathcal{E}_{\text{EFM}}/\partial t=\boldsymbol{J}_{e}\cdot\boldsymbol{E}=0$. When the pressure gradient enters the the Eulerian element, however, the presence of the diamagnetic drift yields a positive $\boldsymbol{J}_{e}\cdot\boldsymbol{E}=W_{\text{diam}}$. At the same time, there is a net outward flux of high-pressure plasma that is being replaced by inflowing low-pressure plasma. In terms of the energy equation, this corresponds to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}>0$, $\nabla\cdot\boldsymbol{Q}>0$, and $\partial\mathcal{E}_{\text{EFM}}/\partial t<0$. This is exactly the situation we find at the island edge around $x/d_{i0}=10$ in Figure 5. We therefore end up with the perhaps counterintuitive conclusion that in the EFM, a positive $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ can lead to a decreasing energy density, due to the work being diverted into energy fluxes. We also note that the $\boldsymbol{J}_{e}\cdot\boldsymbol{E}>0$ region is not the source of the bulk flow, since the entire plasma is drifting, and it would do so even in the absence of the pressure gradient.
@@ 4.2 Extrapolating to the physical mass ratio
One important result from our analysis is that the LGCM can produce significant errors due to the erroneous assumption of $\mu$-conservation when $\kappa<1$. In our simulations, we found regions of low $\kappa$ at the x-points, near the island edges, and inside the islands for $B_{g}=0$ (Figure 4d). Increasing the guide field to $0.2$ increased $\kappa$, particularly inside the islands, but low-$\kappa$ regions still remained. Since the reconnecting current sheet is on electron-scales, increasing the mass-ratio should not affect $\kappa$ near the current sheet center or in the electron diffusion region. Indeed, from spacecraft observations it is well known that electrons demagnetize in these regions (e.g. Burch et al.,, 2016; Torbert et al.,, 2018; Cozzani et al.,, 2019). However, since the magnetic islands grow into ion-scale structures, the associated curvature radius should increase for increasing mass-ratios. We might therefore expect that the LGCM problem of demagnetized electrons at magnetic islands should be less of a concern for physical mass-ratios. However, the islands still grow from electron scales, so this argument would only be valid in the later stages of reconnection. What is more, we know from spacecraft data that electron scale structures with strong gradients can be present in physical reconnection outflows (e.g. Zhou et al.,, 2019; Leonenko et al.,, 2021). It is therefore important to be aware of the fact that the LGCM assumes $\mu$ conservation, and to test that assumption before applying the model.
@@ 5 Conclusions
(i) We find that the energization related to the diamagnetic drift ($W_{\text{diam}}$) is generally, on large scales, the dominant contributor to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ in the EFM also in the case of multiple x-point reconnection (see Figure 2). For large guide fields ($B_{g}=1$), the contribution of parallel electric fields, $W_{\parallel}$, becomes the most important term (albeit still comparable to $W_{\text{diam}}$). The increased importance of $W_{\parallel}$ for larger $B_{g}$ has previously been observed for the LGCM (Dahlin et al.,, 2016). The only clear mass-ratio dependence is found for the Eulerian curvature term ($W_{\text{curv1}}$) in the $B_{g}=0$ case, where it becomes less important as $m_{i}/m_{e}$ is increased from 25 to 100. Locally, other energization terms can contribute significantly to electron energization (see Figure 3).
(ii) We show that the assumption of magnetic moment conservation made by the LGCM is invalid in the low $B_{g}\leq 0.2$ case near the x-points and at the center of the current sheets where the magnetic field is strongly curved. This causes significant errors in the LGCM energization estimate. For $B_{g}=1$, the electrons remain magnetized, and we find a good agreement between the simulation domain integrated LGCM and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. Local deviations from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are produced by the LGCM in all simulations, even when $\mu$ is conserved. This is due to the fact that the LGCM and the EFM describe two fundamentally different quantities. The LGCM describes the evolution of the total energy of the guiding centers currently within the volume element, and it does not capture bulk fluid effects such as the diamagnetic current. In contrast, the EFM describes the evolution of the total electron energy content within the volume and accounts for the total fluid current. The latter includes the effects of energy fluxes through the volume, whereas the former does not. By examining the different terms of the electron energy equation in Figure 5, we find that the energy flux terms are important in balancing $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ on a local level, partly explaining the difference between the two models.
@@
In summary, the EFM describes a complete decomposition of $\boldsymbol{J}_{e}$ into fluid drifts, and it therefore accurately reproduces $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. Local measurements of each constituent energization term, however, need not describe locations where the plasma is either heated or accelerated. In contrast, the LGCM accurately describes the energization of particles if it is used validly, i.e. when $\mu$ is conserved, but it does not describe the local evolution of the bulk fluid energy. A thorough understanding of what one wishes to describe is therefore necessary before one decides to use either of the energization models.

### v2

@@ 3.3 Eulerian vs Lagrangian energization
On system scales, however, the picture is very different. For our choice of boundary conditions, the system is closed in the sense that there is no net particle energy flux into or out of the simulation domain. Therefore, when we integrate the divergence terms of Eq. (4) over the whole simulation, the corresponding surface terms vanish, leaving us with $\boldsymbol{J}_{e}\cdot\boldsymbol{E}\approx\partial\mathcal{E}_{\text{EFM}}/\partial t$, as seen in Figure 5i. This is why the domain-integrated LGCM can accurately reproduce $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ when $\mu$-conservation applies (Figure 2e). In that case, Eq. (2) describes the evolution of the total energy of all guiding centers, and since they are confined to the simulation domain, this is the same thing as the evolution of the total electron energy in the simulation domain. So, when we look at a closed system in its entirety, indeed $\int d^{3}x(d\mathcal{E}_{\text{LGCM}}/dt)=\int d^{3}x(\partial\mathcal{E}_{\text{EFM}}/\partial t)=\int d^{3}x\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. The difference between the domain-integrated LGCM and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ observed in the $B_{g}=0$ case (Figure 2a) can thus be attributed to the lack of $\mu$ conservation due to the small $\kappa$ (Figure 4d). The fact that the LGCM equation is locally invalid in the $B_{g}=0$ case explains why the domain-integrated LGCM both overestimates and underestimates $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ in the $B_{g}=0$ case (Figure 2a), and subsequently why the time-integrated energization more accurately reproduces $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$.
Until this point, we have followed TenBarge et al., (2024) and uncritically assumed that, since $d\mathcal{E}_{\text{LGCM}}/dt$ accurately describes the energization of individual guiding centers, it must equal the work done on the electrons as quantified by $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. In other words, since we can write Eq. (2) as the scalar product of a net current density $\boldsymbol{J}_{e,\text{LGCM}}$ and $\boldsymbol{E}$, $\boldsymbol{J}_{e,\text{LGCM}}$ should be equal to the bulk fluid current density $\boldsymbol{J}_{e}$. This, however, is evidently not the case. As already mentioned by TenBarge et al., (2024), the guiding center model is a particle model that is gyroperiod integrated, and it consequently does not describe certain bulk fluid properties such as the diamagnetic drift and the corresponding currents. Therefore, in general $\boldsymbol{J}_{e,\text{LGCM}}\neq\boldsymbol{J}_{e}$, since $\boldsymbol{J}_{e}$ describes the total electron current. In hindsight, it is therefore quite clear that we should never have expected $d\mathcal{E}_{\text{LGCM}}/dt$ to accurately model $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ to begin with, even when $\mu$ is conserved.
@@ 4.1 Interpretation of EFM energization terms
To illustrate how the above scenario affects the local energy evolution of the plasma, we consider an Eulerian volume element that is initially in the high-pressure region. At the start the uniform plasma is $\boldsymbol{E}\times\boldsymbol{B}$-drifting across the volume element, resulting in $\nabla\cdot\boldsymbol{Q}=\partial\mathcal{E}_{\text{EFM}}/\partial t=\boldsymbol{J}_{e}\cdot\boldsymbol{E}=0$. When the pressure gradient enters the Eulerian element, however, the presence of the diamagnetic drift yields a positive $\boldsymbol{J}_{e}\cdot\boldsymbol{E}=W_{\text{diam}}$. At the same time, there is a net outward flux of high-pressure plasma that is being replaced by inflowing low-pressure plasma. In terms of the energy equation, this corresponds to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}>0$, $\nabla\cdot\boldsymbol{Q}>0$, and $\partial\mathcal{E}_{\text{EFM}}/\partial t<0$. This is exactly the situation we find at the island edge around $x/d_{i0}=10$ in Figure 5. We therefore end up with the perhaps counterintuitive conclusion that in the EFM, a positive $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ can lead to a decreasing energy density, due to the work being diverted into energy fluxes. We also note that the $\boldsymbol{J}_{e}\cdot\boldsymbol{E}>0$ region is not the source of the bulk flow, since the entire plasma is drifting, and it would do so even in the absence of the pressure gradient.
@@ 4.2 Extrapolating to the physical mass ratio
One important result from our analysis is that the LGCM can produce significant errors due to the erroneous assumption of $\mu$-conservation when $\kappa<1$. In our simulations, we found regions of low $\kappa$ at the x-points, near the island edges, and inside the islands for $B_{g}=0$ (Figure 4d). Increasing the guide field to $0.2$ increased $\kappa$, particularly inside the islands, but low-$\kappa$ regions still remained. Since the reconnecting current sheet is on electron-scales, increasing the mass-ratio should not affect $\kappa$ near the current sheet center or in the electron diffusion region. Indeed, from spacecraft observations it is well known that electrons demagnetize in these regions (e.g. Burch et al.,, 2016; Torbert et al.,, 2018; Cozzani et al.,, 2019). However, since the magnetic islands grow into ion-scale structures, the associated curvature radius should increase for increasing mass-ratios. We might therefore expect that the LGCM problem of demagnetized electrons at magnetic islands should be less of a concern for physical mass-ratios. However, the islands still grow from electron scales, so this argument would only be valid in the later stages of reconnection. How well the electron magnetic moment is conserved may also depend on several other parameters like the plasma beta and the current sheet type (Harris versus force-free). These parameters can therefore affect the validity of the LGCM. In any case, we know from spacecraft data that electron scale structures with strong gradients can be present in physical reconnection outflows in geospace (e.g. Zhou et al.,, 2019; Leonenko et al.,, 2021). It is therefore important to be aware of the fact that the LGCM assumes $\mu$ conservation, and to test that assumption before applying the model.
@@ 5 Conclusions
(i) We find that the energization related to the diamagnetic drift ($W_{\text{diam}}$) is generally, on large scales, the dominant contributor to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ in the EFM (see Figure 2). For large guide fields ($B_{g}=1$), the contribution of parallel electric fields, $W_{\parallel}$, becomes the most important term (albeit still comparable to $W_{\text{diam}}$). The increased importance of $W_{\parallel}$ for larger $B_{g}$ has previously been observed for the LGCM (Dahlin et al.,, 2016). The only clear mass-ratio dependence is found for the Eulerian curvature term ($W_{\text{curv1}}$) in the $B_{g}=0$ case, where it becomes less important as $m_{i}/m_{e}$ is increased from 25 to 100. Locally, other energization terms can contribute significantly to electron energization (see Figure 3).
(ii) We show that the assumption of magnetic moment conservation made by the LGCM is invalid in the low $B_{g}\leq 0.2$ case near the x-points and at the center of the current sheets where the magnetic field is strongly curved. This causes significant errors in the LGCM energization estimate. For $B_{g}=1$, the electrons remain magnetized, and we find a good agreement between the simulation domain integrated LGCM and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. Local deviations from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are produced by the LGCM in all simulations, even when $\mu$ is conserved. This is due to the fact that the LGCM and the EFM describe two fundamentally different quantities. The LGCM describes the evolution of the total energy of the guiding centers that are currently within a given volume element. In contrast, the EFM describes the evolution of the total electron energy content within the volume element. The latter includes the effects of energy fluxes through the volume element, whereas the former does not. This has the consequence that the two models can both be correct, while simultaneously giving seemingly contradictory local energization rates if there is a net energy flux into or out of the volume element under consideration. By examining the different terms of the electron energy equation in Figure 5, we find that the energy flux terms are important in balancing $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ on a local level, explaining the difference between the two models when $\mu$ is conserved.
@@
In summary, the EFM describes a complete decomposition of $\boldsymbol{J}_{e}$ into fluid drifts, and it therefore accurately reproduces $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. Local measurements of each constituent energization term, however, need not describe locations where the plasma is either heated or accelerated. In contrast, the LGCM accurately describes the energization of particles if it is used validly, i.e. when $\mu$ is conserved, but it does not describe the local evolution of the bulk fluid energy. Local deviations between the energization measures of the two models, even when $\mu$ is conserved, can be attributed to the contrasting Eulerian and Lagrangian perspectives, and does not mean that one of the models is wrong. A thorough understanding of what one wishes to describe is therefore necessary before one decides to use either of the energization models.
@@ Acknowledgments
Appendix A Generalized adiabaticity estimate
[…]
Figure 6: Difference between $\kappa$ and $\kappa^{*}$ for $B_{g}/B_{\infty}=0$ (left column) and $B_{g}/B_{\infty}=0.2$ (right column). (a) Ratio between the LGCM sum and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. The thick black and grey contours correspond to $\kappa^{*}=1$ and $\kappa=1$, respectively. (b) $\kappa$ calculated using $r_{\text{curv}}$. The thick grey contours correspond to $\kappa=1.$(c) $\kappa^{*}$ calculated using $L_{\nabla\boldsymbol{B}}$. The thick black contours correspond to $\kappa^{*}=1$. (d-f) Same format as panels (a-c). All colormaps are saturated.
In summary, other length and time scales can also affect the adiabaticity of electrons (Stephens et al.,, 2017). However, $\kappa$ is sufficient to show that electrons can be non-adiabatic in both the low and intermediate guide field cases, and that $\kappa<1$ correlates strongly with large deviations between the LGCM sum and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$.
Appendix B Local energization and demagnetization for $B_{g}=1$
In this appendix, we present the data corresponding to Figures 3 and 4 for the $B_{g}=1$ case in Figure 7.
Figure 7: (a-d) Spatial dependence of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the $W$ terms inside magnetic islands for $B_{g}=1$. Same format as Figure 3. (e-h) Deviation between the LGCM sum and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ for $B_{g}=1$. Same format as Figure 4. Note that we have increased the range of the colormap in (h) compared to Figure 4, as $\log_{10}(\kappa)>0.5$ everywhere in this case.
Figure 7c shows that the sum of the LGCM energization terms deviates significantly from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ in the $B_{g}=1$ case. The EFM (Figure 7d), on the other hand, provides an accurate estimate of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. We note that the reversed polarity of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ compared to the $B_{g}=0$ and $B_{g}=0.2$ cases (see Figure 3) is due to the fact that the island in the $B_{g}=1$ case is convecting in the opposite ($-x$) direction.
As shown in Figure 7g, the LGCM tends to provide an energization estimate that is much less than $|\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$, particularly inside the magnetic islands. These deviations cannot be explained by electron demagnetization as quantified by $\kappa$, since $\kappa\gg 1$ throughout the whole simulation domain (Figure 7h).